OpenStudy (anonymous):

A high-jumper with a mass of 65 kg just clears a height of 2.13 m and drops onto a 0.30 m thick landing mat. What average net force does the landing mat exert on the high-jumper if it is compressed by 0.18 m during the landing?

OpenStudy (anonymous):

So far I have done... EP=mgh =65*9.8*2.25 =1433.25 J Therefore EK = 1433.25 J at bottom From EK=1/2mv^2 I found that v=6.64 m s-1 but I'm not sure how to continue to find the force. Would someone please be able to help me?

OpenStudy (xishem):

Work is the change in kinetic energy -- it is also a force applied over a distance, so you should be able to find the change in kinetic energy and relate that to the distance over which the force from the mat was applied.